a) Liala Ltd acquired all the issued shares of Jordan Ltd on 1 January 2015. The following transactions occurred between the two entities:

  •  On 1 June 2016, Liala Ltd sold inventory to Jordan Ltd for $12,000, this inventory previously costed Liala Ltd $10,000. By 30 June 2016, Jordan Ltd had sold 20% of this inventory to other entities for $3,000. The other 80% was all sold to external entities by 30 June 2017 for $13,000.

  •  During the 2016–17 period, Jordan Ltd sold inventory to Liala Ltd for $6,000, this being at cost plus 20% mark-up. Of this inventory, 20 % remained on hand in Liala Ltd at 30 June 2017. The tax rate is 30%.

    Required:

(i) Prepare the consolidation worksheet entries for Liala Ltd at 30 June 2017 in relation to the

intragroup transfers of inventory. (4 marks)
(ii) Compute the amount of cost of goods sold to be reported in the consolidated income

statement for 2017 relating to the relevant intra-group sales. (2 marks)

b) On 1 July 2016, Liala ltd sold an item of plant to Jordan Ltd Ltd for $150,000 when its carrying value in Liala Ltd book was $200,000 (costs $300,000, accumulated depreciation $100,000). This plant has a remaining useful life of five (5) years form the date of sale. The group measures its property plants and equipment using a costs model. Tax rate is 30 percent.

Required:
Prepare the necessary journal entries in 30 June 2017 to eliminate the intra-group transfer of equipment. (4 marks)
Answer #1

Answer :

(a)

(i) Preparation of  the consolidation worksheet entries for Liala Ltd.

Entries on 30 June 2016 :

Date Particulars Debit Credit
30-Jun-16 Sales $12,000
    Cost of Goods Sold $12,000
(To eliminate intra entity sales)

Calculation of unrealized Gross profit in inventory as on 30 June 2016 :

Profit percentage = (2000/12000) x 1001 = 16.67%

80% remaining as on 2016 = 12000 x 80% = 9,600

Unrealized profit = 9600 x 16.67% = 1,600

Date Particulars Debit Credit
30-Jun-16 Cost of Goods Sold $1,600
To Inventory $1,600
(To eliminate intra entity gross profit in ending inventory)

Entries on 30 June 2017 :

No elimination of inter company profit is needed because all of the inter company profit has been realized through resale of the inventory to the external party.

Calculation of unrealized profit:

20% remained = 6000 x 20% = $1,200

Unrealized Profit = 1,200 x (20/120) = $200

Date Particulars Debit Credit
30-Jun-17 Retained Earnings - Bottom $200
To Cost of Goods Sold $200
(To eliminate unrealized profit)

(ii) Calculation of Cost of Goods Sold to be reported in consolidated Income Statement for 2017 relating to the relevant intra-group sales:

Cost of Goods Sold of Upstream = $9,600

Cost of Goods sold downstream = $6,000

Unrealized Profit = $200

Cost of Goods sold in consolidated Income Statement = $9,600 + 6,000 - 200 = $15,400

(b)

Date Particulars Debit Credit
30-Jun-17 Plant $50,000
Loss on sale of Plant   $50,000
(To eliminate effects of intra entity transfer of plant)

At the year end Liala ltd records a depreciation expense of 200000/5) = $40,000

Consolidated depreciation expenses will be ($150,000 / 5) =  $30,000

($40,000 - 30,000) = $10,000 consolidated worksheet adjustment to depreciation expense is necessary